# ASCoT: An Adaptive Self-Correction Chain-of-Thought Method for Late-Stage Fragility in LLMs

### v3

@@ 4.2. Main Results
Figure 5. Performance comparison of ASCoT with varying maximum length constraints. We adjust the maximum length budget when evaluating our method on MATH-500.
We also provide the results of Qwen2.5‐Instruct series models as Table 2. Across the three model scales (3B, 7B, and 14B), ASCoT consistently maintains original performance while achieving substantial CoT compression. For the 3B model, ASCoT at 0.9 compression incurs only a 0.5 point drop in accuracy, while reducing token usage by 16% (314.87 $\rightarrow$ 262.31). As the compression ratio decreases to 0.5, the accuracy degrades from 83.5% to 74.9%, but the token count falls sharply to 170.87(reducing 45.7%). Similar trends hold for the 7B and 14B variants: at Ratio = 0.9, both scales suffer less than a 0.5 point accuracy drop (for the 14B model, there is even a slight improvement in performance), yet token counts shrink by about 15%. Even under aggressive compression (ratio = 0.5), the 14B model retains more than 91.6% accuracy, representing only a 1.5 point reduction, while halving the original token budget. The detailed experimental results can be found in the appendix.
@@
Table 2. Experimental results of the Qwen2.5-Instruct series evaluated on GSM8K. We also provide Accuracy, Tokens and ActRatio.
@@
3B Original 83.7(0.0↓) 314.87
ASCoT 1.0 83.5(0.2↓) 318.32 1.01
0.9 83.2(0.5↓) 262.31 0.83
0.8 82.0(1.7↓) 250.75 0.79
0.7 80.1(3.6↓) 233.15 0.74
0.6 77.6(6.1↓) 199.44 0.63
0.5 74.9(8.8↓) 170.87 0.54
7B Original 91.4(0.0↓) 297.83
ASCoT 1.0 91.8(0.4↑) 294.74 0.99
0.9 91.3(0.1↓) 251.71 0.85
0.8 90.4(1.0↓) 234.95 0.79
0.7 90.3(1.1↓) 216.73 0.73
0.6 88.2(3.2↓) 178.28 0.60
0.5 86.3(5.1↓) 150.98 0.51
14B Original 93.1(0.0↓) 313.11
ASCoT 1.0 93.3(0.2↑) 312.30 0.99
0.9 93.6(0.5↑) 269.76 0.86
0.8 93.5(0.4↑) 246.57 0.79
0.7 93.7(0.6↑) 221.17 0.71
0.6 92.8(0.3↓) 177.40 0.57
0.5 91.6(1.5↓) 158.21 0.51
Figure 4 demonstrates that ASCoT consistently achieves substantial CoT compression across all tested Qwen2.5-Instruct series models (3B, 7B, and 14B). For the 3B model, a moderate compression that reduces token usage by nearly 17% (from 314.87 to 262.31 tokens) incurs only a negligible 0.5-point drop in accuracy. This trend is further amplified in the larger models. With the 7B variant, a 15% token reduction results in a minimal 0.1-point accuracy drop, and a nearly 50% reduction in tokens still maintains accuracy above 86%. Notably, the 14B model showcases ASCoT’s exceptional ability to enhance efficiency. Even when halving the original token budget (a 50% reduction to 158.21 tokens), the accuracy drop is a mere 1.5 percentage points. These results validate ASCoT’s effectiveness in striking a favorable balance between computational efficiency and reasoning fidelity, demonstrating its robustness and scalability across different model sizes.
4.3. Error Propagation via Controlled Injection
To quantitatively investigate the dynamics of error propagation in CoT reasoning, we conducted a comprehensive and systematic error-injection study to rigorously test the “cascading failure” hypothesis. Our experimental setup begins with a baseline of correct CoT solutions for problems from the GSM8K benchmark. For each correct reasoning chain, we first introduce two distinct classes of perturbations into the reasoning steps of the correct CoT sequences. Then, we assess and evaluate how these perturbations affect the overall reasoning process.
The injected errors are not manually added incorrect tokens, but simulated based on real symbolic and numerical mistakes observed during the model’s own generation. By replicating these naturally occurring failure patterns in a controlled manner, we enable systematic analysis of how the position and type of such errors affects the final reasoning outcome.
Numeric errors were introduced to simulate arithmetic miscalculations by slightly altering a single number in the reasoning process (e.g., adding or subtracting a fixed constant value). Symbolic errors were introduced to simulate logical inconsistencies by replacing a primary arithmetic operator with its inverse (e.g. + to -, $\times$ to $\div$). We preemptively set strict constraints to avoid division by zero.
Table 4 reports the impact of single‐step perturbations generated by Qwen2.5-Instruct 7B. For numeric errors, the accuracy drops gradually across positions, with a miscalculation at step 2 causing a decrease of 14.64%, at step 3 a 29.56% drop, and at step 4 a significant reduction of 51.69%. In contrast, symbolic errors induce more severe performance degradation, especially when injected later in the reasoning process. A symbolic error at step 2 results in a large drop of 43.97%,. However, injecting a symbolic error at step 3 (3/3) results in an accuracy reduction of 56.78%. These results emphasize the susceptibility of CoT reasoning to errors, particularly with symbolic mistakes, which cause a more pronounced decline in performance as the reasoning progresses. Table 3 reports initial accuracy before error injection. All the above data are from dataset GSM8K.
@@ 4.3. Error Propagation via Controlled Injection
Table 3. Accuracy for Different Inference Steps in the GSM8K Dataset, using Qwen2.5-Instruct 7B.
@@
9 4 7 57.14
10 11 11 100.00
11 2 2 100.00
4.4. Analysis
Length Budget Analysis To further investigate ASCoT’s efficiency, we analyze its performance under a fixed token budget. As illustrated in Figure 5, we assess our approach under two distinct length-budget settings. In the first, the maximum generation length is scaled by the compression ratio (ranging from 0.5 to 1.0). In the second, we retain the original LLM budget (e.g., 512 tokens in GSM8K and 1024 tokens in MATH) regardless of compression. The results reveal that under the unchanged budget, our method consistently surpasses the baseline: accuracy rises from 42.6% to a peak of 49.4%. Notably, at compression ratios of 0.8 and 0.9, we observe absolute accuracy gains of 0.8 and 0.6 percentage points over the baseline in a smaller tokens, underscoring the potential of compressed CoT reasoning within the same budgetary constraints.
The Late-Stage Fragility of CoT Our findings reveal a phenomenon that challenges the traditional cascading failure model. To illustrate this conceptually, Figure 6 contrasts the model’s ability to recover from an identical type of error when introduced at different stages of the reasoning process. An early-stage error, while potentially disruptive, was often corrected or mitigated by the model’s subsequent reasoning steps. However, an identical error introduced in the final or penultimate step was significantly more likely to go uncorrected, directly leading to a wrong answer.
Our findings do not aim to refute the cascading failure hypothesis entirely, but rather to refine and contextualize it. Previous work suggests that early-stage errors in CoT reasoning tend to propagate forward, compromising subsequent steps and the final answer. However, our analysis reveals that the impact of reasoning errors is not uniform. Errors introduced near the conclusion of reasoning are notably much less likely to be corrected. This indicates that cascading failures do occur, but that late-stage fragility constitutes an additional failure mode not previously emphasized.
Figure 6. A illustration of “Late-Stage Fragility”. The model recovers from an early-stage calculation error, but fails when an identical type of error is introduced in the final step. The tokens that occur related errors are highlighted in orange.
Case Study To provide a clear and detailed qualitative understanding of our method, we present a case study, illustrated in Figure 3. The problem asks for the total number of apples Bob picked over nine days.
As shown in the Initial CoT panel, a standard model correctly structures the reasoning but commits a calculation error in Step 2, incorrectly stating that 125 * 8 = 1,005. This single subtle yet critical flaw propagates directly to the final step, resulting in an incorrect answer of 1,020.
The Evaluation panel demonstrates the intervention of AVM module. The AVM performs a multi-dimensional assessment of the flawed step. It correctly identifies the lack of Logical Validity (0/1) due to the arithmetic error. Simultaneously, it recognizes that the step’s Factual Groundedness and Semantic Clarity & Process Utility are sound. This evaluation allows the AVM to pinpoint the exact failure without discarding the entire step’s valid structure.
Because the step’s risk score exceeds the threshold, the MSCE is activated. As shown in the Final result panel, the MSCE provides the corrected calculation (125 * 8 = 1,000) using a dual-path correction mechanism. This targeted repair allows the model to obtain the correct final answer of 1,015.
Figure 3 illustrates this evaluation process, where an initial CoT with a calculation error is analyzed.
5. Related Work
5.1. LLM Reasoning & Efficient Thinking
The advent of CoT prompting marked a significant leap in the reasoning capabilities of LLMs Marconato et al. (2023); Ding et al. (2024); Liu et al. (2024a). CoT decomposes complex problems into a sequence of manageable intermediate steps, dramatically improving performance on arithmetic, commonsense, and symbolic reasoning tasks Team (2024).
These efficiency-focused methods can be broadly categorized: Prompt-based strategies Xu et al. (2025) aim to guide the model toward more concise reasoning through specific instructions in the prompt itself, such as setting a token budget Han et al. (2024). Inference-time dynamic reasoning Ding et al. (2025) includes early-exit mechanisms that terminate the reasoning process once a certain confidence threshold is met, avoiding unnecessary computation. CoT Distillation Wadhwa et al. (2024); Wang et al. (2023) focuses on transferring the reasoning ability of a large “teacher” model to a smaller “student” model.
Table 4. Error-injection results on GSM8K, using Qwen2.5-Instruct 7B. “All/Err” denotes the total number of reasoning steps and the step which a fault was injected. “Type” indicates the category of the injected error. “Ori Acc” is the model’s accuracy before error-injection. “Fin Acc” is the accuracy after injection. “Accu Drop(abs. points)” quantifies the relative decrease in accuracy.
All/Err Type Ori Acc Fin Acc Accu Drop
@@ 5.1. LLM Reasoning & Efficient Thinking
5.2. CoT Compression and Pruning
The verbosity of CoT Wingate et al. (2022), while beneficial for reasoning, leads to high inference latency and computational costs. This has spurred research into CoT compression and pruning Chuang et al. (2024); Xuefei et al. (2023), which aims to eliminate redundant information without degrading reasoning accuracy. This issue is often termed the “overthinking phenomenon,” Kumar et al. (2025); Chen et al. (2024a) where models generate unnecessarily long chains that can even introduce errors.
To address this bottleneck, researchers have explored various techniques. Token-level pruning Fu et al. (2024); Ma et al. (2023), uses a semantic importance metric to identify and remove less critical tokens from the reasoning chain, training the model to learn shortcuts between essential concepts. Other approaches operate at a more structural level. Prune-on-Logic Liu et al. (2025) transforms a CoT into a logic graph and prunes entire reasoning steps based on a loss importance score. Simpler methods, such as direct truncation of the reasoning chain Wang et al. (2024), have also been shown to be surprisingly effective, suggesting that complete reasoning paths are not always necessary.
5.3. Self-correction Mechanism
The capacity for self-correction is a critical frontier Pan et al. (2023) for enhancing the reliability of LLMs. The general paradigm Kamoi et al. (2024); Yang et al. (2025) involves a multi-stage process: generating an initial response, producing feedback on that response, and then refining it.
Approaches diverge based on the source of the feedback. Intrinsic self-correction Li et al. (2024); Liu et al. (2024c) relies on the model’s own capabilities to critique and improve its output without external tools Cao et al. (2025); Tan et al. (2025). However, a key challenge is that models often fail to identify their own errors, a phenomenon attributed to a lack of error-correction examples in training data Liu et al. (2024b); Anantaprayoon et al. (2025). Extrinsic self-correction incorporates feedback from external sources. This involves using a verifier model to assess the correctness of each step Kim et al. (2025), query external knowledge, or check the consistency across outputs generated from multiple perspectives.
6. Limitations
While ASCoT demonstrates promising performance, we acknowledge that there still exist several limitations that define avenues for future research. The efficacy of the framework is inherently dependent on the self-evaluation capabilities of its base Large Language Model (LLM); if a model cannot fundamentally recognize its own fallacies, the utility of the AVM and MSCE modules is consequently constrained. Furthermore, our current evaluation is focused on numeric and symbolic errors within the specific domain of mathematical reasoning. Therefore, the generalization of ASCoT to other error modalities—such as factual, commonsense, or complex logical errors—and its transferability to diverse long-chain reasoning tasks like code generation, legal analysis, or scientific summarization, remain important open questions requiring further study.
7. Conclusion
In this work, we challenge the conventional “cascading failure” hypothesis in CoT reasoning. Through systematic error-injection experiments, we identify and quantify “Late-Stage Fragility”: a counter-intuitive phenomenon where errors introduced near the end of a reasoning chain are significantly more detrimental than those at the beginning. To counteract this specific vulnerability, we introduce the Adaptive self-correction Chain-of-Thought (ASCoT) method, which is a targeted system that uses an adaptive verification manager to precisely identify and a multi-perspective engine to robustly correct these high-impact, late-stage errors. Our extensive experiments on the GSM8K and MATH benchmarks demonstrate that ASCoT achieves outstanding accuracy, outperforming strong baselines including the computationally expensive self-consistency. These results highlight the effectiveness of ASCoT in addressing late-stage fragility, validating the importance of diagnosing specific failure modes to engineer more effective solutions. This research advocates for a shift from uniform verification to adaptive, position-aware strategies. We hope our investigations pave the way for future advancements in correction mechanisms and enhance the reliability, robustness, and efficiency of LLM reasoning.
@@ Appendix A Appendix
A.1. Experimental Setup
To clearly elucidate the algorithmic pipeline of the ASCoT methodology, we provide its pseudocode. The algorithm commences upon receiving a problem query, which is first processed by the target Large Language Model (LLM) to generate an initial CoT solution, denoted as $CoT_{initial}$. This initial output is subsequently passed to the IRM module for token pruning based on a predefined compression ratio $\gamma$, thereby enhancing efficiency. The compressed CoT then undergoes a step-wise risk assessment by the AVM module. For each step $t_{k}$, the AVM computes a Risk Score $R(t_{k})$, which integrates a multi-dimensional quality assessment $Q(t_{k})$ with a Positional Impact score $I(k)$. If this score exceeds a predefined threshold $\tau$, the step is flagged as high-risk and is routed to the MSCE module for remediation. The MSCE employs a dual-path strategy, encompassing intrinsic reflection and extrinsic independent generation to correct the identified error, yielding the final, refined $CoT_{final}$.
@@ A.1. Experimental Setup
A.2. Error‑Injection Experiments
A.2.1. Error Simulation
@@
A.2.2. Distribution of Results
Table 5 illustrates the proportion of total steps across all problems. We can conclude from the chart that a significant majority of problems are concentrated in the 2-5 step range, with 3-step and 4-step problems being the most prevalent.
@@ A.2.2. Distribution of Results
Table 5. Performance analysis by reasoning step count on the GSM8K dataset, illustrating both the frequency of problems per step count and the corresponding model accuracy.
A.3. Analysis of Late‑Stage Fragility
Early-stage errors often activate the model’s latent self-correction capability, wherein subsequent reasoning steps can reinterpret and compensate for previous mistakes, leading to a correct final answer. This is facilitated by the model maintaining higher semantic entropy and contextual flexibility during the initial stages of reasoning. However, as the reasoning chain progresses, the model increasingly commits to a specific semantic trajectory, gradually narrowing its generative space and becoming less sensitive to alternative interpretations. We refer to this phenomenon as semantic commitment, a state in which the model’s outputs become rigidly anchored to previously generated content. At this point, even minor numerical or symbolic errors near the conclusion are less likely to be detected or corrected, resulting in disproportionately high failure rates.

### v4

@@ 5.2 Main Results
Figure 5: Performance comparison of ASCoT with varying maximum length constraints. We adjust the maximum length budget when evaluating our method on MATH-500.
Figure 4 illustrates ASCoT’s robust scalability across the Qwen2.5-Instruct series. While the 3B model incurs a negligible 0.5-point accuracy drop at 17% compression, this resilience amplifies with model size. The 7B variant maintains over 86% accuracy even with a 50% token reduction. Most notably, the 14B model demonstrates exceptional efficiency retention: halving the token budget results in a mere 1.5% accuracy decline. This trend indicates a positive correlation between model capacity and reasoning redundancy, suggesting that ASCoT effectively exploits the richer semantic structures inherent in larger models to maximize pruning without information loss. This confirm ASCoT’s capability to balance computational efficiency and reasoning fidelity across varying model scales.
5.3 Error Propagation of Controlled Injection
To interrogate the cascading failure hypothesis, we conducted a controlled error-injection study on GSM8K using Qwen2.5-7B. We injected two classes of perturbations mimicking naturally occurring mistakes: Numeric (value deviation) and Symbolic (operator inversion).
Results in Table 4 reveal a Late-Stage Fragility. Contrary to the expectation that early errors propagate most severely, we observe that impact escalates with position. For numeric errors, a miscalculation at the final step (4/4) causes a catastrophic 51.69% accuracy drop, significantly outpacing early-stage errors (14.64% at Step 2). Symbolic errors amplify this trend, with late-stage injections (Step 3/3) causing the most profound degradation (56.78%). These findings suggest that models become increasingly committed to their reasoning trajectory, losing the capacity to self-correct as they approach the solution. Distribution details of the inference steps are provided in Table 3.
@@ 5.3 Error Propagation of Controlled Injection
Table 2: Experimental results of the Qwen2.5-Instruct series evaluated on GSM8K. We report Accuracy (%), Tokens and ActRatio.
@@
3B Original 83.7($0\downarrow$) 314.87
[…]
0.5 74.9($8\downarrow$) 170.87 0.54
7B Original 91.4($0\downarrow$) 297.83
[…]
0.7 90.3($1\downarrow$) 216.73 0.73
[…]
14B Original 93.1($0\downarrow$) 313.11
[…]
5.4 Analysis
Length Budget Analysis To further investigate ASCoT’s efficiency, we analyze its performance under a fixed token budget. As illustrated in Figure 5, we assess our approach under two distinct length-budget settings. In the first, the maximum generation length is scaled by the compression ratio (ranging from 0.5 to 1.0). In the second, we retain the original LLM budget ($e.g.,$ 512 tokens in GSM8K and 1024 tokens in MATH) regardless of compression. The results reveal that under the unchanged budget, our method consistently surpasses the baseline: accuracy rises from 42.6% to a peak of 49.4%. Notably, at compression ratios of 0.8 and 0.9, we observe absolute accuracy gains of 0.8 and 0.6 percentage points over the baseline in a smaller tokens, underscoring the potential of compressed CoT reasoning within the same budget.
The Late-Stage Fragility of CoT Our findings refine the traditional cascading failure model by uncovering Late-Stage Fragility phenomenon. While early-stage errors often trigger latent self-correction mechanisms due to high contextual entropy, identical errors introduced in the final steps are significantly more prone to persist uncorrected (qualitative illustration provided in Appendix C). This suggests that the model forms a rigid semantic commitment as reasoning progresses. Consequently, while cascading failures undeniably exist, Late-Stage Fragility constitutes a distinct, critical failure mode where the model lacks the flexibility to revisit and rectify terminal errors.
Case Study To provide a clear and detailed qualitative understanding of our method, we present a case study. Figure 3 illustrates a qualitative analysis where a standard model correctly structures the reasoning but commits a critical arithmetic error by calculating $125\times 8$ as $1005$. This flaw propagates to an incorrect conclusion. ASCoT effectively addresses this issue as the AVM module flags the step due to low Logical Validity while acknowledging its structural soundness. Crucially, this granular diagnosis prevents the unnecessary discarding of the entire reasoning chain. Instead, the MSCE executes a targeted dual-path repair to synthesize a corrected candidate, rectifying the calculation to $1000$, thereby enabling the model to recover and deduce the correct total of $1015$.
@@ 5.4 Analysis
Table 3: Accuracy for Different Inference Steps in the GSM8K Dataset, using Qwen2.5-Instruct 7B.
@@
9–11 17 20 85.00
5.5 Ablation Study
To rigorously evaluate the contribution of each component within ASCoT, we conducted a multi-level ablation study on the GSM8K dataset. We first validated the necessity of the core modules: IRM and AVM. Subsequently, we performed a fine-grained analysis of the MSCE module to disentangle the efficacy of intrinsic versus extrinsic correction strategies. Our results detailed in Appendix F, confirm that the synergistic integration of position-aware verification and dual-path correction is essential for achieving state-of-the-art performance while minimizing token usage.
Table 4: Error-injection results on GSM8K using Qwen2.5-Instruct 7B. “All/Err” denotes the total steps vs. injected step. “Drop” denotes the accuracy drop.
All/Err Type Ori Acc Fin Acc Drop
@@ 5.5 Ablation Study
6 Conclusion
In this work, we critically re-examine the dynamics of error propagation in CoT reasoning, identifying a counter-intuitive phenomenon termed Late-Stage Fragility. Our systematic analysis reveals that contrary to the prevailing cascading failure assumption, errors introduced in the terminal stages of reasoning impose a disproportionately higher penalty on the final outcome. To address this vulnerability, we introduce ASCoT (Adaptive Self-Correction Chain-of-Thought). By integrating an Intelligent Routing Mechanism with a position-aware Adaptive Verification Manager, ASCoT strategically reallocates computational resources to verify high risk, late-stage steps. Extensive experiments on GSM8K and MATH benchmarks demonstrate that ASCoT achieves a superior trade-off between reasoning fidelity and computational efficiency, outperforming strong baselines including Self-Consistency. Ultimately, our findings advocate for a paradigm shift from uniform verification to adaptive, context-sensitive strategies, establishing a new foundation for reliable and resource-efficient LLM reasoning.
Limitations
While ASCoT effectively mitigates Late-Stage Fragility, its performance naturally correlates with the base model’s verification capabilities. Additionally, our current study primarily targets mathematical reasoning focusing on numeric and symbolic errors. Extending the framework to broader domains ($e.g.,$ code generation) and diverse modalities remains a promising avenue for future exploration.
@@ Appendix A Appendix
Appendix B Algorithmic Details
Algorithm 1 formally outlines the inference pipeline of the proposed ASCoT framework. The process initiates with the Intelligent Routing Mechanism (IRM), which semantically prunes the initial chain $\mathcal{C}_{\text{init}}$ to filter redundant tokens, yielding a compressed sequence $\mathcal{C}_{\text{pruned}}$.
Subsequently, the system iterates through each reasoning step $t_{k}$. At each juncture, the Adaptive Verification Manager (AVM) assesses the step against the accumulated context $S_{\text{ctx}}$, computing a risk score based on both content quality and positional impact. If this score exceeds the predefined threshold $\tau$, the step is flagged as fragile. In such cases, the Multi-Perspective Self-Correction Engine (MSCE) is triggered to generate a rectified step $t_{k}^{\text{corr}}$ via dual-path correction. Conversely, low-risk steps are retained directly. This selective intervention ensures that computational resources are reallocated specifically to fix late-stage vulnerabilities while maintaining overall efficiency.
Algorithm 1 ASCoT: Adaptive Self-Correction Chain-of-Thought
1:Initial CoT $\mathcal{C}_{\text{init}}$; Compression ratio $\gamma$; Risk threshold $\tau$
2:Corrected CoT $\mathcal{C}_{\text{final}}$
3:// Stage 1: Semantic Pruning via IRM
[…]
5:// Stage 2 & 3: Adaptive Verification and Correction
[…]
7:for each step $t_{k}$ in $\mathcal{C}_{\text{pruned}}$ do
8:  // Calculate risk via AVM (Logic, Fact, Clarity, Position)
9:  $R_{k}\leftarrow\mathrm{ComputeRiskScore}(t_{k},S_{\text{ctx}})$
10:  if $R_{k}>\tau$ then
11:   // High risk detected: Trigger MSCE correction
12:   $t_{k}^{\text{corr}}\leftarrow\mathrm{DualPathCorrection}(t_{k},S_{\text{ctx}})$
13:   $\mathrm{Append}(t_{k}^{\text{corr}},\mathcal{C}_{\text{final}})$ $\triangleright$ Use corrected step
14:   $S_{\text{ctx}}\leftarrow\mathrm{UpdateContext}(S_{\text{ctx}},t_{k}^{\text{corr}})$
15:  else
16:   $\mathrm{Append}(t_{k},\mathcal{C}_{\text{final}})$ $\triangleright$ Keep original step
17:   $S_{\text{ctx}}\leftarrow\mathrm{UpdateContext}(S_{\text{ctx}},t_{k})$
18:  end if
19:end for
20:// Final Output Generation
21:return $\mathcal{C}_{\text{final}}$
Appendix C Qualitative Analysis of Late-Stage Fragility
To visually demonstrate the Late-Stage Fragility phenomenon discussed in Section 4.3, we provide a comparative case study in Figure 6. The example contrasts the model’s robust recovery from early-stage perturbations against its failure to correct identical errors in the final reasoning steps.
Figure 6: An illustration of Late-Stage Fragility. The model recovers from an early-stage calculation error, but fails when an identical type of error is introduced in the final step. The tokens that occur related errors are highlighted in orange.
Appendix D Experimental Setup
To clearly elucidate the algorithmic pipeline of the ASCoT methodology, we provide its pseudocode. The algorithm commences upon receiving a problem query, which is first processed by the target Large Language Model (LLM) to generate an initial CoT solution, denoted as $CoT_{initial}$. This initial output is subsequently passed to the IRM module for token pruning based on a predefined compression ratio $\gamma$, thereby enhancing efficiency. The compressed CoT then undergoes a step-wise risk assessment by the AVM module. For each step $t_{k}$, the AVM computes a Risk Score $R(t_{k})$, which integrates a multi-dimensional quality assessment $Q(t_{k})$ with a Positional Impact score $I(k)$. If this score exceeds a predefined threshold $\tau$, the step is flagged as high risk and is routed to the MSCE module for remediation. The MSCE employs a dual-path strategy, encompassing intrinsic reflection and extrinsic independent generation to correct the identified error, yielding the final, refined $CoT_{final}$.
@@ Appendix D Experimental Setup
Appendix E Error‑Injection Experiments
E.1 Error Simulation
@@
Table 5: Performance analysis by reasoning step count on the GSM8K dataset, illustrating both the frequency of problems per step count and the corresponding model accuracy.
E.2 Distribution of Results
Table  5 illustrates the proportion of total steps across all problems. We can conclude from the chart that a significant majority of problems are concentrated in the 2-5 step range, with 3-step and 4-step problems being the most prevalent.
Analysis of Late‑Stage Fragility Early-stage errors often activate the model’s latent self-correction capability, wherein subsequent reasoning steps can reinterpret and compensate for previous mistakes, leading to a correct final answer. This is facilitated by the model maintaining higher semantic entropy and contextual flexibility during the initial stages of reasoning. However, as the reasoning chain progresses, the model increasingly commits to a specific semantic trajectory, gradually narrowing its generative space and becoming less sensitive to alternative interpretations. We refer to this phenomenon as semantic commitment, a state in which the model’s outputs become rigidly anchored to previously generated content. At this point, even minor numerical or symbolic errors near the conclusion are less likely to be detected or corrected, resulting in high failure rates.